The difference between two numbers is 16. Five times the smaller is the same as 8 less than the larger. Find the numbers. I'm having trouble finding the equations.
Answer by stanbon(48535)

The difference between two numbers is 16. Five times the smaller is the same as 8 less than the larger. Find the numbers. I'm having trouble finding the equations.
----------------
Let the numbers be x and y.
Let x be the larger number.
EQUATIONS:
1st: x-y=16
2nd: 5y=x-8
-------------
Solve 1st for x: x=y+16
Substitute in 2nd: 5y=y+16-8
4y=8
y=2
Then x=y+16=18
-------------
larger number is 18; smaller is 2
Cheers,
Stan H.
The larger number is